# BrainECHO: Semantic Brain Signal Decoding through Vector-Quantized Spectrogram Reconstruction for Whisper-Enhanced Text Generation

### v1

@@ Introduction
where $\gamma$ and $\beta_{2}$ are used to scale the latent alignment loss and the commitment loss, respectively. The intermediate representations of the codebook and speech provide additional supervisory signals to guide the generation of Mel spectrograms. We employ L2 loss rather than CLIP loss (Défossez et al. 2023; Yang et al. 2024b) to generate highly restored spectrograms that match Whisper’s input.
Whisper Finetuning
After obtaining the predicted Mel spectrogram, it is fed into the pretrained Whisper-base model to decode tokens. To adequately leverage the pre-trained knowledge embedded in Whisper, we utilize AdaLoRA (Zhang et al. 2023b), as employed in NeuSpeech (Yang et al. 2024a) and MAD (Yang et al. 2024b), to fine-tune its encoder while keeping the remaining parameters frozen. The objective in this final stage is to minimize the cross-entropy loss between the predicted sentence and the ground truth $t$.
@@
Input Method N=1 N=2 N=3 N=4 P R F
Noise NeuSpeech (Yang et al. 2024a) 8.45 1.78 0.43 0 10.26 21.61 13.02 198.31
@@
EEG feature EEG-to-Text (Wang and Ji 2022) 8.82 3.15 1.90 1.44 10.13 21.61 13.12 233.99
EEG NeuSpeech (Yang et al. 2024a) 85.31 84.38 83.98 83.75 82.60 82.73 82.64 16.97
@@
Table 1: Overall comparison of decoding performance on the Brennan dataset. By default, all methods are evaluated without teacher forcing. The metrics with teacher forcing (w/ tf) are further explored. Further results are in the Appendix.
Experiments
Dataset
The Brennan dataset (Brennan and Hale 2019) comprises 49 human EEG recordings, of which 33 remained after screening. Participants passively listened to a 12.4-minute audiobook recording while their EEG signals were recorded. The GWilliams (Gwilliams et al. 2023) dataset contains raw MEG recordings from 27 English speakers who listened to naturalistic stories for 2 hours. More details of the datasets are provided in the Appendix.
Preprocess
@@
All audio is resampled to 16,000 Hz to align with Whisper’s pretraining configuration. To assess the robustness of our proposed method, we employ different approaches to generate samples. For the Brennan dataset, we utilize WhisperX (Bain et al. 2023), a time-accurate speech recognition system, to segment the audio into chunks of up to 12 seconds. For the GWilliams dataset, we split the audio according to the original annotations, resulting in segments no longer than 24 seconds. This process generates a series of EEG/MEG-text-speech pairs.
@@
Dataset Splitting and Validation Strategies
@@
Individual differences and attention levels of subjects can affect EEG signals, making it difficult for models to generalize across subjects and trials. To explore the model’s generalization ability, we design different dataset splitting and validation strategies: random shuffling, session-based, sentence-based, and subject-based splittings. More details are provided in the Appendix. Unless otherwise specified, the Brennan and GWilliams datasets are partitioned by subject-based splittings and random shuffling, respectively, in the following presented results.
Implementation Details
@@
The models are trained on Nvidia 3090 GPUs (24GB). Training on the Brennan and GWilliams datasets take approximately 4 and 24 hours, respectively, using a single GPU. The hyperparameters are set as follows: $\alpha=0.5$, $\beta_{1}=\beta_{2}=0.1$, $\gamma=1$, $N=2048$, $d=256$, and $D=8$. The audio encoder is configured with a downsampling rate of 4. We use vanilla Transformer encoder with 4 layers and 8 heads. All EEG/MEG samples are zero-padded to 2400 in the time dimension. Input spectrograms are padded uniformly to a length of 3000 with -1 following Whisper’s configuration. For the GWilliams dataset, the length of the predicted Mel spectrogram is upsampled by a factor of 2. When generating with Whisper, we set the number of beams to 5 for beam search and apply a repetition penalty of 5.0 with a no-repeat n-gram size of 2. Further details on the training configuration are provided in the Appendix.
Experimental Results
Overall Comparison
@@
We use BLEU (Papineni et al. 2002), ROUGE-1 (Lin 2004) and Word Error Rate (WER) to evaluate decoding performance. BLEU and ROUGE-1 are used to evaluate the quality of text generation, while WER is used to calculate error rate based on edit distance. As shown in Table 1, we compare our model with popularly-referred brain-to-text architectures, EEG-to-Text (Wang and Ji 2022) and NeuSpeech (Yang et al. 2024a). Obviously, our method demonstrates remarkable decoding performance, achieving BLEU-{1, 2, 3, 4} of 89.78, 89.06, 88.74 and 88.55, as well as WER of 11.27 without teacher forcing. The results indicate that BrainECHO generates text highly consistent with the ground truth. Specifically, in terms of BLEU-4, BrainECHO outperforms the previous baseline and current SOTA method by 87.11 (+6049%) and 4.8 (+5.73%) respectively. When using teacher forcing, BrainECHO achieves BLEU-4 of 98.45, which is nearly perfect, highlighting the unrealistic metrics produced by teacher forcing evaluation.
@@
Additionally, when random Gaussian noise is input into BrainECHO, the translation metrics are significantly low, indicating that BrainECHO captures the intrinsic connection between brain signals and text, rather than simply memorizing sentences from the training set. Intuitively, BrainECHO is more resistant to noise than NeuSpeech (Yang et al. 2024a). Notably, the model ideally should not respond to noise, with a WER expected to be 1. Therefore, a high WER ($>$ 1), suggesting the model outputs excessive irrelevant content, is not necessarily a desirable result.
Different Datasets and Splitting Strategies
BLEU-N (%) $\uparrow$
Dataset Split N=1 N=2 N=3 N=4
Brennan Subject 89.78 89.06 88.74 88.55
Sentence 89.24 88.52 88.18 88.01
GWilliams RS 73.35 72.66 72.46 72.42
Session 75.24 74.57 74.34 74.27
Subject 75.05 74.38 74.18 74.14
Sentence 73.58 72.99 72.82 72.79
Table 2: Comparison of decoding performance on different datasets and splits. RS denotes random shuffling.
@@
The decoding metrics of BrainECHO across different datasets and splitting strategies are shown in Table 2. The model demonstrates optimal performance on the Brennan and GWilliams dataset when split by sentences and sessions, respectively. Notably, the performance differences across various splitting strategies are not significant, indicating that BrainECHO is robust and effectively alleviates covariate shift among different subjects or trials without the need for external information (e.g., subject or trial identifiers), provided that all unique sentences are encountered during training. In contrast, the brain module used in (Défossez et al. 2023; Yang et al. 2024b) employs distinct projection matrices for each subject to mitigate individual differences, yet it cannot be generalized to unseen subjects directly.
Examples of Generated Sentences
Some samples generated from different methods are shown in Table 4. These examples indicate that BrainECHO can produce sentences that closely match the original text, even when the reference is long and complex. Remarkably, even without the final fine-tuning of Whisper, BrainECHO still generates results highly relevant to the original text, highlighting the effectiveness of brain-audio latent space alignment (stage 2). In contrast, EEG-to-Text (Wang and Ji 2022) struggles to generate semantically relevant sentences, and NeuSpeech (Yang et al. 2024a) may generate content unrelated to the ground truth when decoding long sentences, which significantly impacts practical applications.
@@
Table 3: Comparison of decoding performance using separate and joint autoencoding on the Brennan dataset. By default, we employ separate autoencoding.
Ablation Study and Analysis
Autoencoding
Generated samples on Brennan (Brennan and Hale 2019)
EEG-to-Text (Wang and Ji 2022) But they were all locked, and when Alice had been all the way down one side and up the other trying every door, she did not care how she was ever to get out again.
NeuSpeech (Yang et al. 2024a) There seemed to be no use in waiting by the little door, so she went back to the table.
EEG-to-Text (Wang and Ji 2022) how she longed to get out of that dark hall and wander about among those beds of bright flowers and those cool fountains, but she did not even get her head through the doorway.
NeuSpeech (Yang et al. 2024a) But they were all locked, and when Alice had been all the way down one side and up the other trying every door, she walked sadly down the middle, wondering how she was ever to get out again.
Generated samples on GWilliams (Gwilliams et al. 2023)
EEG-to-Text (Wang and Ji 2022) It was a long time since I had last seen him in the flesh
NeuSpeech (Yang et al. 2024a) I seen him since high school when I was young, at least before and we were never buddies in any place.
EEG-to-Text (Wang and Ji 2022) He said he had no idea how long it would take him to get back home
NeuSpeech (Yang et al. 2024a) My patience was long gone and I was back in the car. But when I heard that many of you were looking for whatever it was, but what about this?
Table 4: Comparison of decoding sentences generated by different methods, where bold and underline respectively indicates an exact match and a similar match between prediction and ground truth. All methods use the same generation configuration. w/o ft means decoding by inputing the predicted Mel spectrogram into Whisper directly without fine-tuning in the last stage. Additional examples are provided in the Appendix.
We compare the decoding performance when the autoencoding task (stage 1) is applied separately to the Mel spectrograms from individual datasets versus both datasets combined. The results, presented in Table 3, indicate that joint autoencoding results in either stable or slightly improved metrics (except for BLEU-3) compared to separate autoencoding when splitting Brennan by subject. Additionally, all metrics improve when splitting by sentence. This suggests that incorporating Mel spectrograms from other datasets during autoencoding enhances the model’s ability to extract richer discrete speech representations, thereby enhancing its generalizability.
Prediction of Mel Spectrogram
Figure 3: Predicted Mel spectrograms with different window segments on the Brennan dataset. The sentence in this figure is:” that she’d never before seen a rabbit with either a waistcoat pocket or a watch to take out of it, and burning with curiosity, she ran across the field after it, and fortunately.”
@@
Window Length n (s) N=1 N=2 N=3 N=4
12 89.78 89.06 88.74 88.55
6 50.45 45.13 42.54 40.72
4 15.94 5.94 2.60 1.24
Table 5: Decoding comparison for various window lengths in 12-sec Mel spectrogram on Brennan dataset.
Examples of Mel spectrograms predicted by BrainECHO are presented in the Appendix. Since reconstructing a 12-second Mel spectrogram from brain signals in a single pass can significantly impact training efficiency and memory usage, we opt to divide the raw signals into n-second segments to balance decoding accuracy and resource consumption. The corresponding Mel spectrogram for each segment is predicted and then concatenated to form a complete one. We experiment with window lengths of n = {4, 6, 12}, with results shown in Table 5. As the segmentation granularity increases, translation performance gradually deteriorates, indicating that predicting the entire Mel spectrogram at once is more effective than segmenting. The latter method disrupts sentence semantics and introduces discontinuities at segment boundaries. Predicted spectrograms are compared in Figure 3, where clear segmentation artifacts are evident with the splitting method. Notably, if the predicted Mel spectrogram lacks silence at the end, the resulting image appears blurry and smooth, an issue that becomes more pronounced as the window length decreases. This suggests that the model constructs a more robust discrete representation space when using complete Mel spectrograms with silent portions.
Downsampling Ratio
To assess the impact of the downsampling ratio $r$, we evaluate BrainECHO’s performance at $r$ values of 2, 4, 8, and 16, while holding other hyperparameters constant. Assuming each pixel in the spectrogram is represented by 8 bits, the corresponding reductions in bit usage are approximately 2.9, 11.6, 46.5, and 186.1, respectively. As illustrated in Figure 4, increasing $r$ exacerbates information loss, making accurate reconstruction of Mel spectrograms for sentence decoding more challenging. Interestingly, the decoding performance at $r$ = 2 is not as strong as at $r$ = 4, indicating that while a larger feature map enhances reconstruction quality, it may also introduce redundant, translation-irrelevant information, thereby complicating the fine-tuning of Whisper. Therefore, selecting a moderate $r$ is essential to optimize latent representation capacity.
@@
Figure 4: Changes of BLEU-1 and Mel spectrogram reconstruction loss in stage 2 with different downsampling ratio.
Three Training Stages
Training Stage BLEU-N (%) $\uparrow$
Au Al F N=1 N=2 N=3 N=4
89.78 89.06 88.74 88.55
87.13 86.29 85.92 85.74
87.63 86.87 86.54 86.38
39.64 34.49 31.07 28.32
Table 6: Ablation study of training stages on the Brennan dataset. The stages labeled Au, Al, and F correspond to Mel spectrogram autoencoding, brain-audio latent space alignment, and Whisper fine-tuning, respectively.
To verify the effectiveness of our proposed three-stage training, we incrementally remove each stage and observe the corresponding changes in performance. As presented in Table 6, when the autoencoding stage is removed–where the quantizer and audio decoder are randomly initialized–BLEU-4 drops to 85.74 (-3.17%). Further removal of the brain-audio alignment stage leads to an abnormal increase in BLEU, highlighting the challenge of directly constructing a representation space from the brain signals to the Mel spectrogram. However, by pre-warming a discrete representation space, the reconstruction quality and stability are enhanced. The most significant decline in performance occurs when the fine-tuning phase is omitted. Notably, even without fine-tuning, BrainECHO achieves impressive performance based solely on the predicted Mel spectrogram, suggesting that excessive fine-tuning of Whisper is unnecessary, as it could result in overfitting.
Conclusion

### v2

@@ 3.2.2 Brain-Audio Latent Space Alignment
where $\gamma$ and $\beta_{2}$ are used to scale the latent alignment loss and the commitment loss, respectively. The intermediate representations of the codebook and speech provide additional supervisory signals to guide the generation of Mel spectrograms. We employ L2 loss rather than CLIP loss (Défossez et al., 2023; Yang et al., 2024c) to generate highly restored spectrograms that match Whisper’s input.
3.2.3 Whisper Finetuning
After obtaining the predicted Mel spectrogram, it is fed into the pretrained Whisper-base [^1] https://huggingface.co/openai/whisper-base.en model to decode tokens. Guided by both the need for computational efficiency and the proven success of this method in related work (Yang et al., 2024b, c), we utilize AdaLoRA (Zhang et al., 2023) to fine-tune its encoder while keeping the remaining parameters frozen. The objective is to minimize the cross-entropy loss between the predicted sentence and the ground truth $t$. While it is feasible to integrate the previous stages and this stage into one stage for end-to-end training, we adopt a three-stage framework for decoupled representation learning and training cost reduction. More discussion is presented in Appendix B.
@@ 3.2.3 Whisper Finetuning
Split Input Method N=1 N=2 N=3 N=4 P R F
Subject Noise NeuSpeech Yang et al. (2024b) 8.45 1.78 0.43 0 10.26 21.61 13.02 198.31
@@
EEG feature EEG-to-Text Wang and Ji (2022) 8.82 3.15 1.90 1.44 10.13 21.61 13.12 233.99
EEG NeuSpeech Yang et al. (2024b) 85.31 84.38 83.98 83.75 82.60 82.73 82.64 16.97
EEG MAD (Yang et al., 2024c) 80.34 79.10 78.46 78.15 81.00 90.76 83.79 42.14
@@
Sentence EEG BrainECHO 89.24 88.52 88.18 88.01 85.56 85.78 85.63 12.34
Table 1: Overall comparison of decoding performance on the Brennan dataset.
BLEU-N (%) $\uparrow$ ROUGE-1 (%$\uparrow$ WER (%) $\downarrow$
Split Input Method N=1 N=2 N=3 N=4 P R F
Random MEG feature EEG-to-Text Wang and Ji (2022) 9.21 2.13 0.57 0.14 9.74 10.73 11.38 118.25
Shuffling MEG NeuSpeech Yang et al. (2024b) 50.49 46.85 44.42 42.55 46.39 52.48 47.10 71.17
MEG NeuSpeech (Original results) 60.3 55.26 51.24 47.78 60.88 59.76 58.73 56.63
MEG MAD Yang et al. (2024c) 3.93 0.42 0 0 8.98 6.85 7.26 105.33
MEG BrainECHO 73.35 72.66 72.46 72.42 69.66 70.12 69.73 31.44
Session MEG NeuSpeech Yang et al. (2024b) 53.16 - - - - - - -
MEG BrainECHO 75.24 74.57 74.34 74.27 72.94 72.84 72.78 29.59
Subject MEG BrainECHO 75.05 74.38 74.18 74.14 71.83 72.02 71.72 29.80
Table 2: Overall comparison of decoding performance on the GWilliams dataset.
4 Experiments
4.1 Dataset
The Brennan dataset (Brennan and Hale, 2019) comprises 49 human EEG recordings, of which 33 remained after screening. Participants passively listened to a 12.4-minute audiobook recording while their EEG signals were recorded. The GWilliams (Gwilliams et al., 2023) dataset contains raw MEG recordings from 27 English speakers who listened to naturalistic stories for 2 hours. More details are provided in Appendix C.
4.2 Preprocessing
@@
All audio is resampled to 16,000 Hz to align with Whisper’s pretraining configuration. To assess the robustness of our proposed method, we employ different approaches to generate samples. For the Brennan dataset, we utilize WhisperX (Bain et al., 2023), a time-accurate speech recognition system, to segment the audio into chunks of up to 12 seconds. For the GWilliams dataset, we split the audio according to the original annotations, resulting in segments no longer than 24 seconds. This process generates a series of EEG/MEG-text-speech pairs.
@@
4.3 Dataset Splitting and Validation Strategies
@@
Individual differences and attention levels of subjects can affect EEG signals, making it difficult for models to generalize across subjects and trials. To explore the model’s generalization ability, we design different dataset splitting and validation strategies: random shuffling, session-based, sentence-based, and subject-based splittings. More details about the splitting strategies are provided in Appendix D. We ensure that the test data are completely separate from the training data. However, we must note that our data split is done in pairs, i.e., "Semantic Audio – Brain Signal evoked by the Semantic Audio." This means that the same sentence, evoking the same brain signal in a specific trial, will not appear in both the training and testing stages. Unless otherwise specified, the Brennan and GWilliams datasets are partitioned by subject-based splittings and random shuffling, respectively, in the following results.
4.4 Implementation Details
@@
The models are trained on Nvidia 3090 GPUs (24GB). Training on the Brennan and GWilliams datasets take approximately 4 and 24 hours, respectively, using a single GPU. The hyperparameters are set as follows: $\alpha=0.5$, $\beta_{1}=\beta_{2}=0.1$, $\gamma=1$, $N=2048$, $d=256$, and $D=8$. The audio encoder is configured with a downsampling rate of 4. We use a vanilla Transformer encoder with 4 layers and 8 heads. All EEG/MEG samples are zero-padded to 2400 in the time dimension. Input spectrograms are padded uniformly to a length of 3000 with -1 following Whisper’s configuration. For the GWilliams dataset, the length of the predicted Mel spectrogram is upsampled by a factor of 2. When generating with Whisper, we set the number of beams to 5 for beam search and apply a repetition penalty of 5.0 with a no-repeat n-gram size of 2. Further details on the training configuration are provided in Appendix E.
4.5 Comparative Study
We use BLEU (Papineni et al., 2002), ROUGE-1 (Lin, 2004), and Word Error Rate (WER) to evaluate decoding performance. BLEU and ROUGE-1 assess the quality of text generation, while WER calculates error rates based on edit distance.
4.5.1 Benchmarking SOTA Methods on the Brennan Dataset
@@
We compare our model with popularly-referred brain-to-text architectures, i.e., EEG-to-Text (Wang and Ji, 2022), NeuSpeech (Yang et al., 2024b) and MAD (Yang et al., 2024c). NeuSpeech Yang et al. (2024b), the previous SOTA model for MEG-to-text translation, serves as the baseline for comparison. MAD (Yang et al., 2024c) introduces brain-audio alignment on the basis of NeuSpeech. To ensure a fair comparison, we replicated these frameworks based on our data split settings, using the same training and test data as input. As shown in Table 1, our method demonstrates remarkable decoding performance, achieving BLEU-{1, 2, 3, 4} of 89.78, 89.06, 88.74 and 88.55, as well as WER of 11.27 without teacher forcing. The results indicate that BrainECHO generates text highly consistent with the ground truth. Specifically, in terms of BLEU-4, BrainECHO outperforms the previous baseline and current SOTA method by 87.11 (+6049%) and 4.8 (+5.73%) respectively. When using teacher forcing, BrainECHO achieves BLEU-4 of 98.45, which is nearly perfect, highlighting the unrealistic metrics produced by teacher forcing evaluation.
@@
Additionally, since BrainECHO is a generative model, it always produces some output, even with noise, which occasionally matches a few words and gets non-zero BLEU scores. However, the BLEU-4 score of zero shows that matching four consecutive words is unlikely. The noise results are still far from the EEG/MEG results, indicating that BrainECHO captures the intrinsic connection between brain signals and text, rather than simply memorizing sentences from the training set. Intuitively, BrainECHO is more resistant to noise than NeuSpeech (Yang et al., 2024b). Notably, the model ideally should not respond to noise, with a WER expected to be 1. Therefore, a high WER ($>$ 1), suggesting the model outputs excessive irrelevant content, is not necessarily a desirable result.
4.5.2 Benchmarking SOTA Methods on the GWilliams Dataset
Evaluation metrics on the GWilliams dataset across various splitting strategies are presented in Table 2. When using random shuffling, BrainECHO achieves a BLEU-4 score of 72.42, outperforming NeuSpeech by 24.64 points (+51.57%). Furthermore, with session-based division, BrainECHO achieves a BLEU-1 score of 75.24, exceeding NeuSpeech by 22.08 points (+41.53%). These results indicate that BrainECHO can generate text that closely matches the ground truth. Additionally, the results we reproduced on MAD are unsatisfactory on both datasets, especially on GWilliams, indicating that optimizing the CLIP loss between neural signals and audio representations is particularly challenging when the input signal is long (the original experimental setup in MAD used only a 4-second time length). Some examples of generated sentences are presented in Appendix F.
@@ 4.5.2 Benchmarking SOTA Methods on the GWilliams Dataset
Additionally, the performances across various splitting strategies are presented. BrainECHO demonstrates optimal performance on the GWilliams dataset when split by sessions. In particular, the performance differences are not significant, indicating that BrainECHO is robust and effectively alleviates covariate shift among different subjects or trials without the need for external information (e.g., subject or trial identifiers), provided that all unique sentences are encountered during training. In contrast, the brain module used in (Défossez et al., 2023; Yang et al., 2024c) employs distinct projection matrices for each subject to mitigate individual differences, yet it cannot be generalized to unseen subjects directly. More discussion about the rationality of splitting strategies is provided in Appendix H.
Figure 3: Predicted Mel spectrograms on the Brennan dataset (Left: Ground truth Mel spectrogram; Right: Reconstructed Mel spectrogram).
4.5.3 The Reconstructed Mel Spectrograms
Figure 3 shows samples of Mel spectrograms reconstructed from brain signals in the Brennan dataset. The corresponding results for the GWilliams dataset are provided in Appendix F. These samples demonstrate that BrainECHO can produce Mel spectrograms that are largely consistent with the ground truth. Notably, the model effectively restores fine details and accurately predicts the intervals and silent segments in the spectrograms. These results highlight the model’s expressive and predictive capabilities, as it can extract Mel spectrograms from brain signal segments exceeding 20 seconds–a feat not achieved by previous methods.
4.6 Ablation and Hyperparameter Study
4.6.1 Ablation Study on Three-Stage Training
Training Stage BLEU-N (%) $\uparrow$
Au Al F N=1 N=2 N=3 N=4
89.78 89.06 88.74 88.55
87.13 86.29 85.92 85.74
87.63 86.87 86.54 86.38
39.64 34.49 31.07 28.32
Table 3: Ablation study of training stages on the Brennan dataset. The stages labeled Au, Al, and F correspond to Mel autoencoding, brain-audio alignment, and Whisper fine-tuning, respectively.
To verify the effectiveness of our proposed three-stage training, we incrementally remove each stage and observe the corresponding changes in performance. As presented in Table 3, when the autoencoding stage is removed, BLEU-4 drops to 85.74 (-3.17%). Note that in this case, the alignment loss between brain signals and Mel spectrograms in the latent space is removed, while the commitment loss and the reconstruction loss of Mel spectrograms are retained. Further removal of the brain-audio alignment stage means eliminating the reconstruction loss as well. At this point, the model is trained end-to-end. This leads to an abnormal increase in BLEU, highlighting the challenge of directly constructing a representation space from the brain signals to the Mel spectrogram. However, by pre-warming a discrete representation space, the reconstruction quality and stability are enhanced. In the above two cases, the quantizer and audio decoder are randomly initialized and trainable due to the removal of the autoencoding stage. Without fine-tuning in the final stage—i.e., feeding the predicted Mel spectrograms directly into Whisper—the performance is suboptimal. This indicates that the brain-audio alignment is imperfect, and Whisper’s recognition results may deviate from the ground truth. Thus, the fine-tuning stage is crucial to bridge these gaps and improve overall performance.
@@ 4.6.1 Ablation Study on Three-Stage Training
Table 4: Comparison of decoding performance using separate and joint autoencoding (Separate: Autoencoding trained individually on Brennan and GWilliam datasets; Joint: Autoencoding trained on the combined Brennan and GWilliam datasets).
@@
N=1 N=2 N=3 N=4
w/ quantizer 89.78 89.06 88.74 88.55
w/o quantizer 86.46 85.57 85.15 84.90
Table 5: Ablation study of quantizer.
@@
Figure 4: Changes of BLEU-1 and Mel spectrogram reconstruction loss with different downsampling ratios.
4.6.2 Impact of Data Input Strategy in the Autoencoding Stage
This experiment allows us to analyze whether joint training in the autoencoding stage (Stage 1) enhances the model’s ability to learn richer and more generalized representations, thereby improving downstream performance in the following stages. Specifically, we compare two approaches: (1) separate autoencoding, where the model is trained individually on the Mel spectrograms from the Brennan and GWilliam datasets, and (2) joint autoencoding, where the Mel spectrograms from both datasets are combined for training. Following Stage 1, the model proceeds to Stage 2 and Stage 3, which are performed separately on the training sets of Brennan and GWilliam to evaluate the generalizability and dataset-specific performance. The results of the Brennan dataset presented in Table 4 show that, overall, joint autoencoding leads to either stable or slightly improved metrics. However, the improvement is marginal, suggesting that the pre-training datasets need to exhibit high correspondence with the downstream EEG/MEG signals to significantly benefit the decoding framework.
4.6.3 Hyperparameter Analysis of Audio Encoder Module
To assess the impact of the downsampling ratio $r$, we evaluate BrainECHO’s performance at $r$ values of 2, 4, 8, and 16, while holding other hyperparameters constant. Assuming each pixel in the spectrogram is represented by 8 bits, the corresponding reductions in bit usage are approximately 2.9, 11.6, 46.5, and 186.1, respectively. As illustrated in Figure 4, increasing $r$ exacerbates information loss, making accurate reconstruction of Mel spectrograms for sentence decoding more challenging. Interestingly, the decoding performance at $r$ = 2 is not as strong as at $r$ = 4, indicating that while a larger feature map enhances reconstruction quality, it may also introduce translation-irrelevant information, thereby complicating the fine-tuning of Whisper. Therefore, selecting a moderate $r$ is essential to optimize latent representation capacity.
4.6.4 Role of the Discrete Encoding Module
As shown in Table 5, removing the quantizer, i.e., using continuous representation instead of discrete representation, results in a performance decline across all metrics compared to the version with a quantizer. This indicates that discretization can enhance the model’s generalization ability by reducing session-specific noise and facilitating the learning of subject-invariant features.
5 Conclusion